# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 5. Lipschitz continuity of the solution map
In this section, we provide the proof of Theorem 1.4. Namely, we show that provided $f$ satisfies (1.14) in addition to Assumption 1.1, then the solution map is Lipschitz continuous on bounded sets of $\mathbb{E}({\mathbf{R}}^{d})$.
@@
It remains to control $\nabla\mathcal{N}(\psi_{1})-\nabla\mathcal{N}(\psi_{2})$ in $N^{0}([0,T]\times{\mathbf{R}}^{d})$. To that end, we recall that $\nabla\mathcal{N}(\psi_{i})$ can be decomposed by means of the functions $G_{\infty}(\psi_{i}),G_{q}(\psi_{i})$ defined in (2.38). One has that
@@
$\leq\||G_{\infty}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{d}))}+\||G_{q}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}$
$+\left\||G_{\infty}(\psi_{1})-G_{\infty}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}+\left\||G_{q}(\psi_{1})-G_{q}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}.$
@@
$|G_{\infty}(\psi_{1})|\leq C,\quad|G_{q}(\psi_{1})|\leq C(1+|\psi_{1}|^{2\alpha}).$
Further, (1.14) yields that $G_{\infty},G_{q}$ are locally Lipschitz, namely,
(5.4) $\left|G_{\infty}(\psi_{1})-G_{\infty}(\psi_{2})\right|$ $\leq C\left||\psi_{1}|-|\psi_{2}|\right|,$
$\left|G_{q}(\psi_{1})-G_{q}(\psi_{2})\right|$ $\leq C\left(1+|\psi_{1}|^{2\beta}+|\psi_{2}|^{2\beta}\right)\left||\psi_{1}|-|\psi_{2}|\right|,$
wit $\beta=\max\{0,\alpha-\frac{1}{2}\}$. As $|\psi_{i}|\geq 1$ on the support of $G_{q}(\psi_{i})$, we may assume in the following that $\beta\geq 1$.
@@
Case 1: $d=2$: Let the admissible pair $(q_{1},r_{1}))=(\frac{2(\alpha+1)}{\alpha},2(\alpha+1))$, see also (3.1). To bound the first line on the right hand side of (5.3), we observe that
$\||G_{\infty}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{L^{1}([0,T];L^{2}({\mathbf{R}}^{2}))}\leq CT\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{2}))},$
@@
$\||G_{q}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi_{1})^{2\alpha}\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{2}))}.$
@@
$\left\||G_{\infty}(\psi_{1})-G_{\infty}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq C\left\||\psi_{1}|-|\psi_{2}||\nabla\psi_{2}|\right\|_{L^{\frac{4}{3}}([0,T];L^{\frac{4}{3}}({\mathbf{R}}^{2}))}$
@@
$\left\||G_{q}(\psi_{1})-G_{q}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}$
$\leq C\left\|\left(1+|\psi_{1,q}|^{2\beta}+|\psi_{2,q}|^{2\beta}\right)\left||\psi_{1}|-|\psi_{2}|\right|\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}$
$\leq\left(T^{\frac{1}{2}}\|\nabla\psi\|_{L^{4}L^{4}}+T^{\frac{1}{3}}\left(\||\psi_{1,q}|^{2\beta}+|\psi_{1,q}|^{2\beta}\|_{L_{t}^{\infty}L_{x}^{6}}\right)\|\nabla\psi\|_{L_{t}^{3}L_{x}^{6}}\right)\left\||\psi_{1}|-|\psi_{2}|\right\|_{L_{t}^{\infty}L_{x}^{2}}$
@@
where we used the Strichartz estimates (2.29), (2.25) and (3.4) in the last inequality. Combining the above estimates, we obtain that there exists $T_{1}=T_{1}(M)>0$ sufficiently small so that
@@
Acknowledgments
P.A. and P.M. acknowledge partial support from INdAM-GNAMPA. P.A. is partially supported by PRIN project 20204NT8W4 ”Nonlinear evolution PDEs, fluid dynamics and transport equations: theoretical foundations and applications”. L.E.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.

### v4

@@ 5. Lipschitz continuity of the solution map
In this section, we provide the proof of Theorem 1.4. Namely, we show that provided $f$ satisfies (1.14) in addition to Assumption 1.1, the solution map is Lipschitz continuous on bounded sets of $\mathbb{E}({\mathbf{R}}^{d})$.
@@
It remains to control $\nabla\mathcal{N}(\psi_{1})-\nabla\mathcal{N}(\psi_{2})$ in $N^{0}([0,T]\times{\mathbf{R}}^{d})$. To that end, we recall that $\nabla\mathcal{N}(\psi_{i})$ can be decomposed by means of the functions $G_{\mathrm{bd}}(\psi_{i}),G_{\mathrm{int}}(\psi_{i})$ defined in (2.37). One has that
@@
$\leq\||G_{\mathrm{bd}}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{d}))}+\||G_{\mathrm{int}}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}$
$+\left\||G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}+\left\||G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{d})}.$
@@
$|G_{\mathrm{bd}}(\psi_{1})|\leq C,\quad\text{\quad}|G_{\mathrm{int}}(\psi_{1})|\leq C(1+|\psi_{1}|^{2\alpha}).$
Further, (1.14) yields that $G_{\mathrm{bd}}$ and $G_{\mathrm{int}}$ are locally Lipschitz, namely,
(5.4) $\left|G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi_{2})\right|$ $\leq C\left||\psi_{1}|-|\psi_{2}|\right|,$
$\left|G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi_{2})\right|$ $\leq C\left(1+|\psi_{1}|^{2\beta}+|\psi_{2}|^{2\beta}\right)\left||\psi_{1}|-|\psi_{2}|\right|,$
wit $\beta=\max\{0,\alpha-\frac{1}{2}\}$. As $|\psi_{i}|\geq 1$ on the support of $G_{\mathrm{int}}(\psi_{i})$, we may assume in the following that $\beta\geq 1$.
@@
Case 1: $d=2$: Consider the admissible pair $(q_{1},r_{1}))=(\frac{2(\alpha+1)}{\alpha},2(\alpha+1))$, see also (3.1). To bound the first line on the right hand side of (5.3), we observe that
$\||G_{\mathrm{bd}}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{L^{1}([0,T];L^{2}({\mathbf{R}}^{2}))}\leq CT\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{2}))},$
@@
$\||G_{\mathrm{int}}(\psi_{1})||\nabla\psi_{1}-\nabla\psi_{2}|\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi_{1})^{2\alpha}\|\nabla\psi_{1}-\nabla\psi_{2}\|_{L^{\infty}([0,T];L^{2}({\mathbf{R}}^{2}))}.$
@@
$\left\||G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq C\left\||\psi_{1}|-|\psi_{2}||\nabla\psi_{2}|\right\|_{L^{\frac{4}{3}}([0,T];L^{\frac{4}{3}}({\mathbf{R}}^{2}))}$
@@
$\left\||G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi_{2})||\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}$
$\leq C\left\|\left(1+|\psi_{1,\mathrm{int}}|^{2\beta}+|\psi_{2,\mathrm{int}}|^{2\beta}\right)\left||\psi_{1}|-|\psi_{2}|\right|\nabla\psi_{2}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}$
$\leq\left(T^{\frac{1}{2}}\|\nabla\psi\|_{L^{4}L^{4}}+T^{\frac{1}{3}}\left(\||\psi_{1,\mathrm{int}}|^{2\beta}+|\psi_{1,\mathrm{int}}|^{2\beta}\|_{L_{t}^{\infty}L_{x}^{6}}\right)\|\nabla\psi\|_{L_{t}^{3}L_{x}^{6}}\right)\left\||\psi_{1}|-|\psi_{2}|\right\|_{L_{t}^{\infty}L_{x}^{2}}$
@@
where we used the Strichartz estimates (2.28), (2.24) and (3.4) in the last inequality. Combining the above estimates, we obtain that there exists a sufficiently small $T_{1}=T_{1}(M)>0$ so that
@@
Acknowledgements
P.A. and P.M. acknowledge partial support from INdAM-GNAMPA. P.A. is partially supported by PRIN project 20204NT8W4 ”Nonlinear evolution PDEs, fluid dynamics and transport equations: theoretical foundations and applications” and by the Italian Ministry of University and Research (MUR) through the Excellence Department Project awarded to GSSI, CUP D13C22003740001. L.E.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.